the perimeter of a rectangle is 42 ft. The length of the rectangle is 3 ft less than twice the width.Find the length and the width of the rectangle
Answer by mananth(16946)

width = w
length = 2w-3
...
2*(l+w)=perimeter
2*(w+2w-3)=42
2*(3w-3)=42
6w-6=42
+6 +6
6w=42+6
6w=48
6w/6=48/6
w=8 the width.
..
length = 2w-3
length = 2*8-3
16-3
=13
..
length =13feet, width =8feet
..
m.ananth@hotmail.ca